OpenStudy (anonymous):

Which of the following equations is of a parabola with a vertex at (0, 6)?

OpenStudy (anonymous):

y = (x - 6)^2 y = (x + 6)^2 y = x^2 - 6 y = x^2 + 6

OpenStudy (anonymous):

The equation of a parabola with the vertex \((h,k)\) is: \[ y-k=a(x-h)^2 \]

OpenStudy (anonymous):

thanks